# AI-SearchPlanner: Modular Agentic Search via Pareto-Optimal Multi-Objective Reinforcement Learning

### v3

@@ 1 Introduction
Recently, reinforcement learning (RL) has emerged as a promising paradigm for improving the reasoning capabilities of LLMs. Models such as OpenAI-o1111https://openai.com/zh-Hans-CN/index/learning-to-reason-with-llms/ and DeepSeek-R1 Guo et al. (2025) employ RL techniques (e.g., PPO Schulman et al. (2017), GRPO Shao et al. (2024)) to enhance logical reasoning and problem-solving through iterative experiential learning. Research like Search-R1 Jin et al. (2025) further highlights the potential of RL by leveraging question-answering rewards in multi-turn reasoning to train LLMs for search engine interactions. This approach yields significant performance improvements, demonstrating RL’s effectiveness in enhancing LLM-based search utilization. Although RL-based search agents have demonstrated superior performance in complex reasoning tasks, existing approaches typically rely on a single LLM to handle both search planning and QA tasks in an end-to-end manner, which poses significant challenges in optimizing both capabilities simultaneously. In particular, in practical scenarios, sophisticated AI search systems such as Baidu [^2] https://chat.baidu.com/ and Tencent Yuanbao [^3] https://yuanbao.tencent.com/chat/ commonly utilize a large frozen LLM (e.g., GPT-4 Achiam et al. (2023), DeepSeek-R1 Guo et al. (2025)) to ensure high-quality QA. Thus, a more effective and efficient solution is to employ a small, trainable LLM dedicated to search planning, thereby improving QA performance while maintaining low computational latency.
@@ 2 Related Work
Although large language models (LLMs) Vaswani et al. (2017); Achiam et al. (2023); Zhao et al. (2023) exhibit remarkable general question-answering capabilities, they still lack the ability to leverage external domain-specific knowledge to solve complex reasoning problems. To mitigate these limitations, search engines are commonly employed to augment LLMs with external information. There are two primary approaches to integrating search engines with LLMs: (1) retrieval-augmented generation (RAG) Lewis et al. (2020); Gao et al. (2023); Mei et al. (2025) and (2) treating the search engine as a tool within an agent framework Schick et al. (2023); Jin et al. (2025); Qu et al. (2025); Song et al. (2025); Wu et al. (2025a); Li et al. (2025a). Retrieval-Augmented Generation (RAG) Lewis et al. (2020); Gao et al. (2023); Mei et al. (2025) enhances LLMs by integrating retrieved external knowledge into the input, thereby providing access to up-to-date or domain-specific information. However, most existing RAG systems Ma et al. (2023); Jiang et al. (2023b; a) follow rigid retrieval-generation pipelines, which lack adaptive planning mechanisms for complex reasoning. As a result, their performance remains heavily dependent on retrieval accuracy, limiting their robustness in dynamic or knowledge-intensive scenarios.
@@
With the advancement of the general capabilities of LLMs, recent research Qu et al. (2025); Trivedi et al. (2022a); Yao et al. (2023); Schick et al. (2023) has explored integrating search engines as tools to enhance problem-solving. Specially, prompt-based methods Trivedi et al. (2022a); Yao et al. (2023) and self-supervised fine-tuning (SFT) approaches Schick et al. (2023) have been proposed to enable LLMs to autonomously determine when and how to invoke search engines during reasoning. For instance, IRCoT Trivedi et al. (2022a) and ReAct Yao et al. (2023) utilize iterative prompting strategies to interleave reasoning with search engine calls, while Toolformer Schick et al. (2023) leverages supervised fine-tuning to improve search-related capabilities. However, prompt-based methods often exhibit limited generalization, especially for complex reasoning tasks that lie outside the pretraining distribution of LLMs. SFT-based methods often depend on large-scale, high-quality annotated trajectories of search-reasoning interactions, which limits their scalability. Moreover, recent studies suggest that SFT-based approaches tend to overfit to memorized reasoning paths, further constraining their generalization to novel or out-of-distribution scenarios. Recent studies Jin et al. (2025); Song et al. (2025); Wu et al. (2025a); Li et al. (2025a; c) demonstrate that reinforcement learning (RL) can enable LLMs to develop advanced reasoning capabilities using only outcome-based rewards. For example, Search-R1-like approaches Jin et al. (2025); Song et al. (2025) have investigated the potential of RL methods such as PPO Schulman et al. (2017) and GRPO Shao et al. (2024) in agentic search scenarios. By employing outcome-based reward mechanisms and end-to-end training, these methods enable LLMs to jointly optimize multi-turn search planning and answer generation, achieving strong QA performance. A related concurrent work is s3 Jiang et al. (2025), which also decouples the planner from the generator. Our approach differs from s3 as follows: 1) Objective: s3 aims to train a high-performing planner with minimal data, while we focuses on industrial applications, optimizing the planner for the best utility-cost trade-off using state-of-the-art LLMs. 2) Framework: s3 provides the generator with key documents from the final reasoning step; we instead supply the entire planning trajectory. 3) Method Design: s3 introduces a Gain Beyond RAG reward and filters out training examples solvable by naive RAG, thereby improving end-to-end QA performance on complex questions. We directly optimizes planning ability with an outcome-process dual reward and incorporates Pareto optimization of utility and cost, significantly reducing cost while maintaining performance. Additionally, another similar line of recent work Xiao et al. (2024); Shao et al. (2025), ReasoningIR, enhance IR models with intrinsic reasoning ability, by using LLM reasoning to generate query-document pairs for training. In contrast, we leverages query–search–document planning trajectories to directly produce high-quality answers.
@@ 3.4 Reinforcement Learning Training
  • Optimization of Search Planner. We optimize the search planning policy using reinforcement learning with a dual reward formulation. Each search planning trajectory comprises an input question, a sequence of retrieved documents, and a termination decision (i.e., call generator to answer). Upon constructing the final context, a frozen generator generates the answer, and the dual reward is computed. In this work, we employ Proximal Policy Optimization (PPO) Schulman et al. (2017). The corresponding PPO objective is defined as:
@@ 4.1 Datasets
To evaluate AI-SearchPlanner, we use seven Wikipedia-based datasets and two Web-based datasets: 1) Wiki-based: NQ Kwiatkowski et al. (2019), TriviaQA Joshi et al. (2017), PopQA Mallen et al. (2022), HotpotQA Yang et al. (2018), 2WikiMultiHopQA Ho et al. (2020), Musique Trivedi et al. (2022b), and Bamboogle Press et al. (2022). Specially, the first three are general QA datasets, and the last four are multi-hop QA datasets. 2) Web-based: WebShaper Tao et al. (2025) and WebWalkerQA Wu et al. (2025b).
@@ 4.2 Baselines
To evaluate the effectiveness of AI SearchPlanner, we compare it with the following baseline methods that do not decouple the planner and generator: 1) Inference without Retrieval: Direct inference and Chain-of-Thought (CoT) reasoning Wei et al. (2022). 2) Inference with Retrieval: Naive Retrieval-Augmented Generation (RAG) Lewis et al. (2020), IRCoT Trivedi et al. (2022a), and Search-o1 Li et al. (2025b). 3) Training-based Methods: Supervised fine-tuning (SFT) Chung et al. (2024) and Search-R1 Jin et al. (2025).
@@ 4.3 Experimental Settings
For training AI-SearchPlanner, We merge the training sets of NQ and HotpotQA to form a unified dataset. Qwen-2.5-7b-Instruct [^4] https://huggingface.co/Qwen/Qwen2.5-7B-Instruct and Qwen3-32b [^5] https://huggingface.co/Qwen/Qwen3-32B were used as the trainable planner and frozen generator, respectively. The maximum thresholds for the number of reasoning turns and sub-queries $M_{t}$ and $M_{q}$ is set to 5 and 10, respectively. The prompts for evaluating process reward and answer accuracy are provided in Appendix B. Since different frozen QA models may produce answers with divergent content and stylistic variations, the Exact Match (EM) metric is inadequate for measuring answer correctness. To address this limitation, we introduce an LLM-based scoring function $Score\left(\cdot\right)$ that evaluates the answer accuracy of model responses based on the question and the ground truth answer. On the Wikipedia-based datasets, we follow the settings in Jin et al. (2025) to conduct the experiments. We use the 2018 Wikipedia dump Karpukhin et al. (2020) as the knowledge source and E5 Wang et al. (2022) as the retriever. The number of retrieved passages is set to 3. The maximum number of reasoning turns is set to 5.
@@ 4.4.3 Training Dynamics
$\alpha$ = 0
0.674
[[1.856]]
0.800
[[1.910]]
0.540
[[1.749]]
0.678
[[2.352]]
0.565
[[2.681]]
0.355
[[2.867]]
0.568
[[2.408]]
0.597
[[2.260]]
$\alpha$ = 0.005
0.669
[[[1.388]]
0.809
[[1.304]]
0.546
[[1.546]]
0.677
[[1.961]]
0.570
[[2.452]]
0.368
[[2.476]]
0.600
[[2.072]]
0.606(+1.51%)
[[1.886]](-16.55%)
$\alpha$ = 0.05
0.671
[[1.135]]
0.798
[[1.157]]
0.543
[[1.248]]
0.657
[[1.858]]
0.535
[[2.313]]
0.346
[[2.596]]
0.640
[[1.976]]
0.599(+0.34%)
[[1.755]](-22.35%)
$\alpha$ = 0.125
0.662
[[1.000]]
0.795
[[1.000]]
0.542
[[1.000]]
0.564
[[1.000]]
0.385
[[1.000]]
0.219
[[1.000]]
0.400
[[1.000]]
0.510(-14.57%)
[[1.000]](-55.75%)
$\alpha$ = 0.25
0.650
[[1.000]]
0.797
[[1.000]]
0.564
[[1.000]]
0.560
[[1.000]]
0.380
[[1.000]]
0.204
[[1.000]]
0.416
[[1.000]]
0.510(-14.57%)
[[1.000]](-55.75%)

### v4

@@ 1 Introduction
Recently, reinforcement learning (RL) has emerged as a promising paradigm for improving the reasoning capabilities of LLMs. Models such as OpenAI-o1111https://openai.com/zh-Hans-CN/index/learning-to-reason-with-llms/ and DeepSeek-R1 guo2025deepseek employ RL techniques (e.g., PPO schulman2017proximal, GRPO shao2024deepseekmath) to enhance logical reasoning and problem-solving through iterative experiential learning. Research like Search-R1 jin2025search further highlights the potential of RL by leveraging question-answering rewards in multi-turn reasoning to train LLMs for search engine interactions. This approach yields significant performance improvements, demonstrating RL’s effectiveness in enhancing LLM-based search utilization. Although RL-based search agents have demonstrated superior performance in complex reasoning tasks, existing approaches typically rely on a single LLM to handle both search planning and QA tasks in an end-to-end manner, which poses significant challenges in optimizing both capabilities simultaneously. In particular, in practical scenarios, sophisticated AI search systems such as Baidu [^2] https://chat.baidu.com/ and Tencent Yuanbao [^3] https://yuanbao.tencent.com/chat/ commonly utilize a large frozen LLM (e.g., GPT-4 achiam2023gpt, DeepSeek-R1 guo2025deepseek) to ensure high-quality QA. Thus, a more effective and efficient solution is to employ a small, trainable LLM dedicated to search planning, thereby improving QA performance while maintaining low computational latency.
@@ 2 Related Work
Although large language models (LLMs) vaswani2017attention; achiam2023gpt; zhao2023survey exhibit remarkable general question-answering capabilities, they still lack the ability to leverage external domain-specific knowledge to solve complex reasoning problems. To mitigate these limitations, search engines are commonly employed to augment LLMs with external information. There are two primary approaches to integrating search engines with LLMs: (1) retrieval-augmented generation (RAG) lewis2020retrieval; gao2023retrieval; mei2025survey and (2) treating the search engine as a tool within an agent framework schick2023toolformer; jin2025search; qu2025tool; song2025r1; wu2025webdancer; li2025websailor. Retrieval-Augmented Generation (RAG) lewis2020retrieval; gao2023retrieval; mei2025survey enhances LLMs by integrating retrieved external knowledge into the input, thereby providing access to up-to-date or domain-specific information. However, most existing RAG systems ma2023query; jiang2023active; jiang2023llmlingua follow rigid retrieval-generation pipelines, which lack adaptive planning mechanisms for complex reasoning. As a result, their performance remains heavily dependent on retrieval accuracy, limiting their robustness in dynamic or knowledge-intensive scenarios.
@@
With the advancement of the general capabilities of LLMs, recent research qu2025tool; trivedi2022interleaving; yao2023react; schick2023toolformer has explored integrating search engines as tools to enhance problem-solving. Specially, prompt-based methods trivedi2022interleaving; yao2023react and self-supervised fine-tuning (SFT) approaches schick2023toolformer have been proposed to enable LLMs to autonomously determine when and how to invoke search engines during reasoning. For instance, IRCoT trivedi2022interleaving and ReAct yao2023react utilize iterative prompting strategies to interleave reasoning with search engine calls, while Toolformer schick2023toolformer leverages supervised fine-tuning to improve search-related capabilities. However, prompt-based methods often exhibit limited generalization, especially for complex reasoning tasks that lie outside the pretraining distribution of LLMs. SFT-based methods often depend on large-scale, high-quality annotated trajectories of search-reasoning interactions, which limits their scalability. Moreover, recent studies suggest that SFT-based approaches tend to overfit to memorized reasoning paths, further constraining their generalization to novel or out-of-distribution scenarios. Recent studies jin2025search; song2025r1; wu2025webdancer; li2025websailor; li2025towards demonstrate that reinforcement learning (RL) can enable LLMs to develop advanced reasoning capabilities using only outcome-based rewards. For example, Search-R1-like approaches jin2025search; song2025r1 have investigated the potential of RL methods such as PPO schulman2017proximal and GRPO shao2024deepseekmath in agentic search scenarios. By employing outcome-based reward mechanisms and end-to-end training, these methods enable LLMs to jointly optimize multi-turn search planning and answer generation, achieving strong QA performance. A related concurrent work is s3 jiang2025s3, which also decouples the planner from the generator. Our approach differs from s3 as follows: 1) Objective: s3 aims to train a high-performing planner with minimal data, while we focuses on industrial applications, optimizing the planner for the best utility-cost trade-off using state-of-the-art LLMs. 2) Framework: s3 provides the generator with key documents from the final reasoning step; we instead supply the entire planning trajectory. 3) Method Design: s3 introduces a Gain Beyond RAG reward and filters out training examples solvable by naive RAG, thereby improving end-to-end QA performance on complex questions. We directly optimizes planning ability with an outcome-process dual reward and incorporates Pareto optimization of utility and cost, significantly reducing cost while maintaining performance. Additionally, another similar line of recent work xiao2024rar; shao2025reasonir, ReasoningIR, enhance IR models with intrinsic reasoning ability, by using LLM reasoning to generate query-document pairs for training. In contrast, we leverages query–search–document planning trajectories to directly produce high-quality answers.
@@ 3.4 Reinforcement Learning Training
  • Optimization of Search Planner. We optimize the search planning policy using reinforcement learning with a dual reward formulation. Each search planning trajectory comprises an input question, a sequence of retrieved documents, and a termination decision (i.e., call generator to answer). Upon constructing the final context, a frozen generator generates the answer, and the dual reward is computed. In this work, we employ Proximal Policy Optimization (PPO) schulman2017proximal. The corresponding PPO objective is defined as:
@@ 4.1 Datasets
To evaluate AI-SearchPlanner, we use seven Wikipedia-based datasets and two Web-based datasets: 1) Wiki-based: NQ kwiatkowski2019natural, TriviaQA joshi2017triviaqa, PopQA mallen2022not, HotpotQA yang2018hotpotqa, 2WikiMultiHopQA ho2020constructing, Musique trivedi2022musique, and Bamboogle press2022measuring. Specially, the first three are general QA datasets, and the last four are multi-hop QA datasets. 2) Web-based: WebShaper tao2025webshaper and WebWalkerQA wu2025webwalker.
@@ 4.2 Baselines
To evaluate the effectiveness of AI SearchPlanner, we compare it with the following baseline methods that do not decouple the planner and generator: 1) Inference without Retrieval: Direct inference and Chain-of-Thought (CoT) reasoning wei2022chain. 2) Inference with Retrieval: Naive Retrieval-Augmented Generation (RAG) lewis2020retrieval, IRCoT trivedi2022interleaving, and Search-o1 li2025search. 3) Training-based Methods: Supervised fine-tuning (SFT) chung2024scaling and Search-R1 jin2025search.
@@ 4.3 Experimental Settings
For training AI-SearchPlanner, We merge the training sets of NQ and HotpotQA to form a unified dataset. Qwen-2.5-7b-Instruct [^4] https://huggingface.co/Qwen/Qwen2.5-7B-Instruct and Qwen3-32b [^5] https://huggingface.co/Qwen/Qwen3-32B were used as the trainable planner and frozen generator, respectively. The maximum thresholds for the number of reasoning turns and sub-queries $M_{t}$ and $M_{q}$ is set to 5 and 10, respectively. The prompts for evaluating process reward and answer accuracy are provided in Appendix B. Since different frozen QA models may produce answers with divergent content and stylistic variations, the Exact Match (EM) metric is inadequate for measuring answer correctness. To address this limitation, we introduce an LLM-based scoring function $Score\left(\cdot\right)$ that evaluates the answer accuracy of model responses based on the question and the ground truth answer. On the Wikipedia-based datasets, we follow the settings in jin2025search to conduct the experiments. We use the 2018 Wikipedia dump karpukhin2020dense as the knowledge source and E5 wang2022text as the retriever. The number of retrieved passages is set to 3. The maximum number of reasoning turns is set to 5.
@@ 4.4.3 Training Dynamics
$\alpha$ = 0 0.674 [[1.856]] 0.800 [[1.910]] 0.540 [[1.749]] 0.678 [[2.352]] 0.565 [[2.681]] 0.355 [[2.867]] 0.568 [[2.408]] 0.597 [[2.260]]
$\alpha$ = 0.005 0.669 [[[1.388]] 0.809 [[1.304]] 0.546 [[1.546]] 0.677 [[1.961]] 0.570 [[2.452]] 0.368 [[2.476]] 0.600 [[2.072]] 0.606(+1.51%) [[1.886]](-16.55%)
$\alpha$ = 0.05 0.671 [[1.135]] 0.798 [[1.157]] 0.543 [[1.248]] 0.657 [[1.858]] 0.535 [[2.313]] 0.346 [[2.596]] 0.640 [[1.976]] 0.599(+0.34%) [[1.755]](-22.35%)
$\alpha$ = 0.125 0.662 [[1.000]] 0.795 [[1.000]] 0.542 [[1.000]] 0.564 [[1.000]] 0.385 [[1.000]] 0.219 [[1.000]] 0.400 [[1.000]] 0.510(-14.57%) [[1.000]](-55.75%)
$\alpha$ = 0.25 0.650 [[1.000]] 0.797 [[1.000]] 0.564 [[1.000]] 0.560 [[1.000]] 0.380 [[1.000]] 0.204 [[1.000]] 0.416 [[1.000]] 0.510(-14.57%) [[1.000]](-55.75%)